OpenStudy (anonymous):

What will the temperature be in degrees Fahrenheit when it is 80° Celsius outside? (Recall the formula F = (9C 4072-01-01-03-00_files/i0190000.jpg5) + 32. a. 176° Fahrenheit c. 150.4° Fahrenheit b. 76.0° Fahrenheit d. 201.6° Fahrenheit @FriedRice

OpenStudy (anonymous):

lol formula all messed up with a apicture

OpenStudy (anonymous):

f=(9c/5)

OpenStudy (anonymous):

F = ((9)(80)) / 5 F = 720 / 5 = ?

OpenStudy (anonymous):

144

OpenStudy (anonymous):

144 + 32 = 176 degrees Fahrenheit

OpenStudy (anonymous):

so its A

OpenStudy (anonymous):

do you get it?